# On Geometric Structures for Policy Parameterization in Continuous Control

### v2

@@ Abstract
Standard stochastic policies for continuous control often rely on ad-hoc boundary-enforcing transformations (e.g., tanh) which can distort the underlying optimization landscape and introduce gradient pathologies. While alternative parameterizations on the unit manifold (e.g., directional distributions) are theoretically appealing, their computational complexity (often requiring special functions or rejection sampling) has limited their practical use. We propose a novel, computationally efficient action generation paradigm that preserves the structural benefits of operating on a unit manifold. Our method decomposes the action into a deterministic directional vector and a learnable concentration scalar, enabling efficient interpolation between the target direction and uniform noise on the unit manifold. This design can reduce policy head parameters by nearly 50% (from $2d$ to $d+1$) and maintains a simple $O(d)$ sampling complexity, avoiding costly sampling procedures. Empirically, our method matches or exceeds state-of-the-art methods on standard continuous control benchmarks, with significant improvements (e.g., +37.6% and +112%) on high-dimensional locomotion tasks. Ablation studies confirm that both the unit-norm normalization and the adaptive concentration mechanism are essential to the method’s success. These findings suggest that robust, efficient control can be achieved by explicitly respecting the structure of bounded action spaces, rather than relying on complex, unbounded distributions. Code is available in supplementary materials.
@@ 1 Introduction
  • We provide comprehensive empirical evaluation across six MuJoCo benchmarks, achieving the best performance on 4 out of 6 tasks, especially in high-dimensional control, validating that geometric consistency outweighs distributional sophistication.
@@ 2.2 Beyond Gaussian: Alternative Distributions
The most principled alternative emerged from directional statistics (Sinii et al., 2024). vMF distributions, operating directly on the unit sphere, elegantly address boundary constraints through their geometric formulation. Recent work (Scott et al., 2021) demonstrated vMF policies could match or exceed Gaussian performance while providing theoretical advantages. However, vMF’s practical adoption faces significant hurdles (Banerjee et al., 2005): sampling requires rejection methods with acceptance rates as low as 0.1 for high concentrations, likelihood computation involves modified Bessel functions $I_{v}(\kappa)$ prone to numerical overflow for large $\kappa$, and the concentration parameter lacks intuitive interpretation for practitioners (Zaghloul & Johnson, 2025). These challenges have confined vMF policies largely to theoretical investigations rather than practical deployment.
@@ 2.3 Geometric Perspectives in RL
Parallel to distributional innovations, a geometric understanding of RL has emerged (Hu et al., 2022). Hyperbolic RL leverages non-Euclidean geometry for hierarchical representations, while work on Riemannian policy optimization extends natural gradients to curved manifolds (Nickel & Kiela, 2017; Wang et al., 2024; Müller & Montúfar, 2024). These approaches highlight how geometric structure can inform algorithm design, yet they typically add complexity rather than simplifying existing methods. Our work inverts this trend, using geometric insights to simplify rather than sophisticate.
The connection between action spaces and geometric structures has appeared implicitly in several contexts (Wang et al., 2025). Quaternion representations for rotational actions in robotics naturally embrace spherical geometry, while circular distributions model periodic actions in locomotion tasks (Zhou et al., 2019). However, these remain specialized solutions rather than general principles. GAC unifies previously scattered geometric insights into a single, practical framework for continuous control—demonstrating that the central question is not which distribution to adopt, but whether explicit distributions are necessary at all.
@@ 3.2 Geometric Action Generation
where $\bm{\xi}\sim\text{Uniform}(\mathbb{S}^{d-1})$ is sampled as normalized Gaussian noise to provide isotropic exploration on the unit sphere, and $r=2.5$ is the action radius (task-dependent). Even with substantial noise contribution (e.g., 30% when $\kappa\approx 1$), actions remain coherent: spherical normalization preserves directionality while preventing magnitude corruption, ensuring stable control throughout training.
@@
Gradient Flow. All GAC operations, including normalization, sigmoid mixing, and linear interpolation, are differentiable and compatible with standard backpropagation, ensuring stable policy optimization.
@@ 3.3 Theoretical Justification
This result is exact and requires no approximation: $w(\kappa)$ directly controls the expected alignment with $\bm{\mu}$, and as $\kappa$ grows large ($w(\kappa)\to 1$), variance shrinks and samples concentrate tightly around $\bm{\mu}$. This provides vMF-like concentration behavior without Bessel function computations. See Appendix A.1 for the proof.
@@ 3.4.1 Exploration Control Mechanism
GAC introduces a learned exploration controller $\kappa(s)$ that adaptively modulates the balance between deterministic actions and stochastic exploration. Unlike SAC’s temperature parameter $\alpha$ which requires manual tuning or scheduling, $\kappa$ learns directly from the value landscape.
In the maximum entropy framework, the actor seeks to maximize expected return while maintaining exploration. For GAC, this objective becomes:
$L_{\text{actor}}(\phi)=\mathbb{E}_{s\sim\mathcal{D}}\left[-\kappa(s)-\min_{i=1,2}Q_{\theta_{i}}(s,\mathbf{a})\right],$ (4)
where $\mathcal{D}$ is the replay buffer, $\phi$ denotes the actor parameters, and $\mathbf{a}$ is generated via GAC’s geometric mechanism in (2) with current state $s$. The term $\kappa(s)$ serves as a learned exploration controller that replaces entropy regularization. Smaller values of $\kappa$ increase the contribution of stochastic noise in the geometric mixing defined in (2), thereby promoting exploration. In contrast, larger values lead to more deterministic actions. This removes the need for temperature tuning in SAC. Unlike traditional entropy-based methods, GAC never computes probability densities. Instead, exploration emerges directly from geometric structure. This design is theoretically justified by directional statistics, where higher concentration naturally corresponds to lower entropy (see Appendix A.4).
The soft Q-function update follows standard SAC with our exploration controller. The target value incorporates the minimum of two Q-networks for stability:
@@ 3.5 Practical Considerations
Parameter Efficiency. GAC requires only $d+1$ outputs (direction vector plus scalar concentration) compared to $2d$ for diagonal Gaussian policies, where parameter count reduces by nearly 50%.
Computational Efficiency. The sampling procedure involves only normalization and linear interpolation, avoiding rejection sampling or special function evaluations. The computational complexity is $O(d)$ per sample, compared to $O(dk)$ for vMF sampling where $k$ is the expected number of rejections (typically $k\in[2,10]$ for high concentrations).
Hyperparameter Selection. Due to the geometry of high-dimensional spheres, a unit vector in $\mathbb{R}^{d}$ has expected per-dimension magnitude $\mathbb{E}[|\mu_{i}|]\approx 1/\sqrt{d}$, which becomes vanishingly small as $d$ increases (e.g., $\approx 0.24$ for $d=17$). Without scaling, such weak actions would be ineffective for control. The radius $r=2.5$ provides principled rescaling: after spherical mixing with typical $w(\kappa)\approx 0.85$, this yields per-dimension actions around 0.60.60.90.9, which fall comfortably within the normalized action bounds [1,1][-1,1] and ensure effective actuation. While this default works robustly across diverse tasks, certain environments benefit from adjusted scaling (e.g., Ant-v4’s multi-leg coordination performs better with $r=1.0$ for finer control).
@@ 4 Experiments
4.1 Experimental Setup
Environments. We evaluate GAC across six standard MuJoCo (Todorov et al., 2012) continuous control benchmarks that collectively test different aspects of control: HalfCheetah-v4 (speed optimization), Ant-v4 (multi-leg coordination), Humanoid-v4 (complex balance), Walker2d-v4 (bipedal locomotion), Hopper-v4 (unstable dynamics), and Pusher-v4 (object manipulation). These environments span action dimensions from 3 to 17 and present diverse challenges from optimization to precise coordination.
@@ 4.2 Main Results
Table 1: Average return over final training steps (1M for all tasks, 500K for Pusher-v4). Bold indicates best performance. GAC achieves competitive performance across diverse control tasks.
@@
High-dimensional Control Excellence. GAC shows particular strength in complex, high-dimensional tasks. On Ant-v4 (8D action space), GAC achieves 5633 $\pm$ 158, significantly outperforming SAC (4094 $\pm$ 1039) by 37.6% and TD3 (3531 $\pm$ 1263) by 59.5%. The large variance in SAC’s performance ($\pm$ 1039) suggests instability in Gaussian policies for multi-leg coordination, while GAC maintains remarkable consistency ($\pm$ 158). On HalfCheetah-v4 (6D), GAC achieves the highest return of 12750 $\pm$ 758, slightly exceeding SAC and outperforming TD3 by 4.4%.
@@
Trade-off Between Stability and Expressiveness. GAC demonstrates consistent learning stability across environments, achieving the lowest variance on Pusher-v4 ($32\pm 0$) and competitive performance on Walker2d, HalfCheetah, Ant and Humanoid. This stability advantage is particularly pronounced in high-dimensional spaces, where GAC’s bounded geometric structure prevents the exploration instabilities inherent to squashed Gaussian policies.
However, its performance on Hopper-v4 ($1952\pm 285$) lags behind TD3 ($2896\pm 749$), and similarly underperforms on Pusher-v4. We attribute this to a fundamental trade-off introduced by spherical normalization: while it ensures consistent gradient flow and exploration stability, it imposes a geometric prior that assumes the optimal actions lie close to a unit hypersphere. This assumption aligns well with high-dimensional locomotion tasks, where the optimal actions often reside on a spherical shell due to norm concentration effects, but may be suboptimal for tasks involving asymmetric or contact-rich dynamics. For instance, Hopper’s single-leg locomotion requires strong, unbalanced thrusts that extend beyond the unit sphere, and Pusher’s manipulation demands induce anisotropic action distributions better captured by elliptical or non-spherical structures. These observations suggest future directions for adaptive or task-conditioned geometric priors that modulate between spherical, elliptical, or unconstrained action manifolds based on task characteristics.
Challenging Environments. Humanoid-v4 (17D action space) proves challenging for all methods. During training, GAC obtains 5823 $\pm$ 121, TD3 achieves 5819 $\pm$ 278, while SAC achieves 5717 $\pm$ 123. However, post-training evaluation of GAC achieves 6591 $\pm$ 53, suggesting that the algorithm learns high-quality policies while maintaining exploration during training (detailed evaluation protocol in supplementary materials). This validates that spherical normalization effectively handles complex action spaces and balances exploration-exploitation trade-offs.
4.3 Ablation Studies
We conduct comprehensive ablation studies to analyze the contribution of each component in GAC. Table 2 and Figure 3 presents results on HalfCheetah-v4 across 5 random seeds, systematically dissecting the effect of target magnitude $r$ and key architectural components.
Target magnitude $r$ critically affects performance. The choice of $r=2.5$ is theoretically motivated by high-dimensional sphere geometry: a unit vector in $\mathbb{R}^{d}$ has expected per-dimension magnitude $\mathbb{E}[|\mu_{i}|]\approx 1/\sqrt{d}$, which becomes vanishingly small as $d$ increases (e.g., $\approx 0.24$ for $d=17$), rendering unscaled actions ineffective for control. Figure 3 empirically validates this insight: reducing $r$ from 2.5 to 1.5 causes catastrophic 43% performance degradation due to limited exploration and actuation. In contrast, increasing $r$ to 3.5 has minimal impact ($4.1\%$), suggesting saturation. The default $r=2.5$, after spherical mixing with typical $w(\kappa)\approx 0.85$, yields per-dimension actions in the range 0.6–0.9. This range falls well within the normalized bounds [1,1][-1,1] and strikes an effective balance for HalfCheetah-v4’s locomotion dynamics.

### v3

# Beyond Distributions: Geometric Action Control for Continuous Reinforcement Learning

@@ Abstract
Gaussian policies have dominated continuous control in deep reinforcement learning (RL), yet they suffer from a fundamental mismatch: their unbounded support requires ad-hoc squashing functions that distort the geometry of bounded action spaces. While von Mises-Fisher (vMF) distributions offer a theoretically grounded alternative on the sphere, their reliance on Bessel functions and rejection sampling hinders practical adoption. We propose Geometric Action Control (GAC), a novel action generation paradigm that preserves the geometric benefits of spherical distributions while simplifying computation. GAC decomposes action generation into a direction vector and a learnable concentration parameter, enabling efficient interpolation between deterministic actions and uniform spherical noise. This design reduces parameter count from $2d$ to $d+1$, and avoids the $O(dk)$ complexity of vMF rejection sampling, achieving simple $O(d)$ operations. Empirically, GAC consistently matches or exceeds state-of-the-art methods across six MuJoCo benchmarks, achieving 37.6% improvement over SAC on Ant-v4 and up to 112% on complex DMControl tasks, demonstrating strong performance across diverse benchmarks. Our ablation studies reveal that both spherical normalization and adaptive concentration control are essential to GAC’s success. These findings suggest that robust and efficient continuous control does not require complex distributions, but a principled respect for the geometry of action spaces.
@@ 1 Introduction
  • We provide comprehensive empirical evaluation across MuJoCo and DMControl benchmarks, demonstrating consistent improvements especially in high-dimensional control, validating that geometric consistency outweighs distributional sophistication.
@@ 2.2 Beyond Gaussian: Alternative Distributions
The most principled alternative emerged from directional statistics (Sinii et al., 2024). vMF distributions, operating directly on the unit sphere, elegantly address boundary constraints through their geometric formulation. Recent work (Scott et al., 2021) demonstrated vMF policies could match or exceed Gaussian performance while providing theoretical advantages. However, vMF’s practical adoption faces significant hurdles (Banerjee et al., 2005): sampling requires rejection methods with acceptance rates as low as 0.1 for high concentrations, likelihood computation involves modified Bessel functions $I_{v}(\kappa)$ prone to numerical overflow for large $\kappa$, and the concentration parameter lacks intuitive interpretation for practitioners (Zaghloul and Johnson, 2025). These challenges have confined vMF policies largely to theoretical investigations rather than practical deployment.
@@ 2.3 Geometric Perspectives in RL
Parallel to distributional innovations, a geometric perspective on RL has gained traction (Hu et al., 2022). Hyperbolic RL and Riemannian policy optimization extend learning to non-Euclidean manifolds (Nickel and Kiela, 2017; Wang et al., 2024; Müller and Montúfar, 2024), offering richer representations but often at the cost of added algorithmic complexity. In contrast, our work leverages geometry to simplify rather than sophisticate. The link between action spaces and geometry also surfaces in domain-specific contexts, such as quaternion-based rotation policies or circular distributions for periodic locomotion (Wang et al., 2025; Zhou et al., 2019), yet these insights remain fragmented. GAC unifies these scattered geometric intuitions into a cohesive framework for continuous control—shifting the focus from choosing the right distribution to questioning whether explicit distributions are necessary at all.
@@ 3.2 Geometric Action Generation
where $\bm{\xi}\sim\text{Uniform}(\mathbb{S}^{d-1})$ is sampled as normalized Gaussian noise to provide isotropic exploration on the unit sphere, and $r$ is a task-dependent scaling parameter (default $r=2.5$, see Sec. 3.5 for details). Even with substantial noise contribution (e.g., 30% when $\kappa\approx 1$), actions remain coherent: spherical normalization preserves directionality while preventing magnitude corruption, ensuring stable control throughout training.
@@ 3.3 Theoretical Justification
This result is exact: $w(\kappa)$ directly controls the expected alignment with $\bm{\mu}$. As $\kappa$ increases ($w(\kappa)!\to!1$), variance vanishes and samples concentrate around $\bm{\mu}$, yielding vMF-like concentration without Bessel function computations. See Appendix A.1 for the proof.
@@ 3.4.1 Exploration Control Mechanism
GAC introduces a learned exploration controller $\kappa(s)$ that adaptively modulates the balance between deterministic actions and stochastic exploration. Unlike SAC’s temperature parameter $\alpha$ which requires manual tuning or scheduling, $\kappa$ learns directly from the value landscape. In the maximum entropy framework, the actor seeks to maximize expected return while maintaining exploration. For GAC, this objective becomes:
$L_{\text{actor}}(\phi)=\mathbb{E}_{s\sim\mathcal{D}}\left[\kappa(s)-\min_{i=1,2}Q_{\theta_{i}}(s,\mathbf{a})\right],$ (4)
where $\mathcal{D}$ is the replay buffer, $\phi$ denotes the actor parameters, and $\mathbf{a}$ is generated via GAC’s geometric mechanism in (2) with current state $s$. The term $\kappa(s)$ serves as a learned exploration controller that replaces entropy regularization. Smaller values of $\kappa$ increase the contribution of stochastic noise in the geometric mixing defined in (2), thereby promoting exploration. In contrast, larger values lead to more deterministic actions. This removes the need for temperature tuning in SAC. Unlike traditional entropy-based methods, GAC never computes probability densities. Instead, exploration emerges directly from geometric structure. This design is theoretically justified by directional statistics, where higher concentration naturally corresponds to lower entropy (see Appendix A.4). The soft Q-function update follows standard SAC with our exploration controller. The target value incorporates the minimum of two Q-networks for stability:
@@ 3.5 Practical Considerations
Parameter Efficiency. The fixed-radius GAC requires only $d+1$ outputs (direction vector plus scalar concentration), compared to $2d$ for diagonal Gaussian policies (a 50% reduction in action head parameters). The adaptive scaling variant (Eq. 6) uses $2d+1$ outputs, comparable to Gaussian policies but with explicit geometric structure that decouples direction, exploration, and magnitude. Computational Efficiency. The sampling procedure involves only normalization and linear interpolation, avoiding rejection sampling or special function evaluations. The computational complexity is $O(d)$ per sample, compared to $O(dk)$ for vMF sampling where $k$ is the expected number of rejections (typically $k\in[2,10]$ for high concentrations).
Hyperparameter Selection. The geometry of high-dimensional spheres implies that a unit vector in $\mathbb{R}^{d}$ has an expected per-dimension magnitude of $\mathbb{E}[|\mu_{i}|]\approx 1/\sqrt{d}$ (e.g., $\approx 0.24$ for $d=17$). Without scaling, these inherently small magnitudes would yield ineffective actions. We introduce a fixed radius $r$ for principled rescaling. With $r=2.5$ and a typical mixing weight $w(\kappa)\approx 0.85$, the resulting per-dimension actions fall in the 0.60.60.90.9 range, well within the normalized bounds [1,1][-1,1] for effective actuation.
While this default $r=2.5$ works robustly across diverse tasks, we find specific environments like Ant-v4 benefit from adjusted scaling ($r=1.0$) for finer multi-leg control. Crucially, GAC exhibits low sensitivity to the exact value of $r$; performance within the range $[1.0,3.5]$ typically varies by less than $10\%$ (see Appendix D.2). This confirms that $r$ acts as a stable geometric scaling factor, not a fragile hyperparameter. For tasks requiring maximal flexibility, we next introduce a learnable per-dimension variant.
Adaptive Magnitude Scaling. To handle environments with asymmetric dynamics or fine-grained coordination needs, we extend GAC with a learnable magnitude vector $\mathbf{r}\in\mathbb{R}^{d}$. This variant replaces the scalar $r$ in Eq. (2):
$\mathbf{a}=\mathbf{r}\odot\text{normalize}\left(w(\kappa)\cdot\bm{\mu}+(1-w(\kappa))\cdot\bm{\xi}\right),$ (6)
where $\odot$ denotes element-wise multiplication. The vector $\mathbf{r}$ is produced by a small network head with a softplus activation to ensure positivity. This adaptive scaling preserves GAC’s core geometric structure while enabling a degree of fine-grained control that exceeds the capabilities of traditional Gaussian policies, which rely solely on variance tuning. Experiments in Section 4.2 validate the effectiveness of this enhanced variant on complex DMControl tasks.
@@ 4 Experiments
Environments. We evaluate GAC on six standard MuJoCo benchmarks: HalfCheetah, Ant, Humanoid, Walker2d, Hopper, and Pusher, with action dimensions ranging from 3 to 17.
@@ 4.1 Main Results
Table 1: Performance on MuJoCo benchmarks with fixed action radius ($r=2.5$ for most tasks, $r=1.0$ for Ant-v4). Bold indicates best performance.
@@
High-Dimensional Control. GAC demonstrates strong performance and robustness across complex, high-dimensional tasks. On Humanoid-v4 (17D), it achieves 5823 $\pm$ 121 during training, matching TD3 (5819 $\pm$ 278) and exceeding SAC (5717 $\pm$ 123). More notably, GAC reaches 6591 $\pm$ 53 in post-training evaluation, indicating its ability to learn high-quality policies while maintaining effective exploration. On Ant-v4 (8D), GAC outperforms SAC by 37.6% (5633 $\pm$ 158 vs. 4094 $\pm$ 1039) and TD3 by 59.5% (3531 $\pm$ 1263), with significantly lower variance—highlighting GAC’s stability in multi-leg coordination. On HalfCheetah-v4 (6D), GAC achieves the highest return of 12750 $\pm$ 758, slightly exceeding SAC and outperforming TD3 by 4.4%. These results collectively validate that spherical normalization and geometric action modeling enable GAC to scale gracefully with action dimensionality, achieving both reliable exploration and consistent policy quality in challenging continuous control settings.
Trade-off Between Stability and Expressiveness. GAC achieves stable learning with low variance across most environments, particularly in high-dimensional tasks where its bounded geometry mitigates the exploration instabilities of squashed Gaussians. However, it underperforms on Hopper-v4 and Pusher-v4, revealing a trade-off: spherical normalization stabilizes training but imposes a geometric prior favoring near-unit-norm actions. While effective for norm-concentrated locomotion tasks, this constraint can limit asymmetric or contact-rich control. These limitations motivate our learnable scaling variant (Eq. 6, Sec. 4.2), which introduces per-dimension adaptivity while preserving GAC’s geometric structure.
Table 2: Performance on DMControl suite. GAC uses the adaptive scaling variant (Eq. 6).
Environment GAC (Ours) SAC TD3 PPO
fish-upright 858 $\pm$ 35 923 $\pm$ 5 866 $\pm$ 39 311 $\pm$ 78
walker-walk 960 $\pm$ 4 956 $\pm$ 10 952 $\pm$ 5 186 $\pm$ 11
walker-run 742 $\pm$ 15 700 $\pm$ 56 651 $\pm$ 87 69 $\pm$ 16
cheetah-run 762 $\pm$ 24 661 $\pm$ 185 753 $\pm$ 27 150 $\pm$ 32
quadruped-walk 925 $\pm$ 17 690 $\pm$ 336 873 $\pm$ 94 131 $\pm$ 18
quadruped-run 638 $\pm$ 75 301 $\pm$ 8 576 $\pm$ 212 119 $\pm$ 22
Figure 3: Learning curves on (a) fish-upright, (b) walker-walk, (c) walker-run, (d) cheetah-run, (e) quadruped-walk, and (f) quadruped-run.
4.2 Generalization to DMControl Suite
We further evaluate GAC’s generalization on six challenging tasks from the DMControl suite (Tassa et al., 2018), including fish-upright, walker-walk, walker-run, cheetah-run, quadruped-walk, and quadruped-run, with more complex contacts, asymmetric coordination, and higher-dimensional action spaces than standard MuJoCo tasks. To accommodate these challenges, we adopt the GAC-Scale variant (Eq. 6), which augments the core geometric structure of GAC with learnable per-dimension magnitude control. All experiments follow the same protocol as described earlier.
As shown in Table 2 and Figure 3, GAC-Scale matches or surpasses SAC on 5 out of 6 tasks, with particularly strong gains on quadruped environments (+34%–+112%). These tasks require coordinated multi-leg movement and diverse actuation patterns, capabilities that are difficult to achieve with fixed-radius or diagonal Gaussian policies. The learned scaling vector $\mathbf{r}$ enables fine-grained, task-adaptive control while preserving the exploration benefits of GAC’s spherical normalization.
Analysis of scale dynamics (Appendix D) shows that $\mathbf{r}$ values stabilize within the theoretically motivated range $[1.5,3.0]$ after initial exploration ($\sim$100K steps), with task-specific convergence profiles. For instance, Walker tasks settle to nearly uniform values around 1.71.72.42.4, closely matching the fixed $r=2.5$ baseline. In contrast, Quadruped tasks exhibit more heterogeneous scaling patterns across dimensions, reflecting the need for anisotropic control across legs and joints.
These results confirm that GAC’s geometric framework generalizes effectively to complex, contact-rich environments. The fixed-radius design is sufficient for symmetric locomotion, while the learnable $\mathbf{r}$ extension introduces critical flexibility in asymmetric scenarios, achieving substantial performance gains without compromising the architectural simplicity or stability of the base method.
@@ 4.2 Generalization to DMControl Suite
4.3 Ablation Studies
We conduct ablations on HalfCheetah-v4 to assess the target magnitude $r$ (Table 3, Figure 4).
Target magnitude $r$ is critical. Geometrically, a unit vector in $\mathbb{R}^{d}$ has expected per-dimension magnitude $\mathbb{E}[|\mu_{i}|]\approx 1/\sqrt{d}$ (e.g., $\sim$0.24 for $d=17$), making unscaled actions too weak for control. This is reflected in our results: reducing $r$ from 2.5 to 1.5 causes a 43% performance drop due to insufficient actuation and exploration, while increasing $r$ to 3.5 yields only a minor decline ($4.1\%$), indicating saturation. With $r=2.5$ and typical mixing weight $w(\kappa)\!\approx\!0.85$, the resulting action amplitudes ($\sim$0.6–0.9) fall neatly within [1,1][-1,1], providing an effective balance for locomotion.